OpenStudy (anonymous):

eleven cards say M A T H E M A T I C S are in a box. what is the probability of selecting m and then another m with replacement of the first one? please help!!!

OpenStudy (anonymous):

help!!!!!!

OpenStudy (anonymous):

4/121

OpenStudy (anonymous):

thank you so much i couldnt understand it!!

OpenStudy (anonymous):

suppose i had two cards, M and A. What is the probability that you would choose M?

OpenStudy (anonymous):

P(X1 = M) = 2/11 P(X2 = M) = 2/11 P(X1 = M AND X2 = M) = P(X1 = M given that X2 = M) . P(X2=M) = 2/11 . 2/11 (since the probabilities don't depend on each other). P(X1 = M AND X2 = M) = 4/121

OpenStudy (anonymous):

Oh, I see I took too long.

OpenStudy (anonymous):

lol.